# Using Machine Learning to Explore Defect Configurations in Cd/Zn-Se/Te Compounds

### v1

@@ Abstract
We developed a framework for predicting the energies and ground state configurations of native point defects, extrinsic dopants and impurities, and defect complexes across zincblende-phase Cd/Zn-Te/Se/S compounds, important for CdTe-based solar cells. This framework, named “DeFecT-FF”, is powered by high-throughput density functional theory (DFT) computations and crystal graph-based machine learning force field (MLFF) models trained on the DFT data. The Cd/Zn-Te/Se/S chemical space is chosen because alloying at Cd or Te sites is a promising avenue to tailor the electronic and defect properties of the CdTe absorber layer to potentially improve solar cell performance. The sheer number of defect configurations achievable when considering all possible singular defects and their combinations, symmetry-breaking operations, and defect charge states, as well as the expense of running large supercell calculations, makes this an ideal problem for developing accurate and widely-applicable force field models. Here, we introduce our datasets of structures and energies from GGA-PBE and HSE06 geometry optimization, including bulk and alloyed supercells with and without defects, and defect-containing interface and grain boundary structures. Datasets were gradually expanded using active learning and accurate MLFF models were trained to predict energies and atomic forces across different charge states. Via accelerated prediction and screening, we identified many new low energy defect configurations and obtained high-fidelity defect formation energy diagrams using HSE06 calculations with spin-orbit coupling. The DeFecT-FF framework has been released publicly as a nanoHUB tool, allowing users to upload any crystallographic information file, generate defects of interest, and compute defect formation energies as a function of Fermi level and chemical potential conditions, thus bypassing expensive DFT calculations.
@@
Md Habibur Rahmana and Arun Mannodi-Kanakkithodia
@@ 1 Introduction
[…]
Shockley-Read-Hall (SRH) recombination is a major limiting factor that reduces the power conversion efficiency and is often associated with grain boundaries, point defects, and dislocations 6. Native point defects and impurities can introduce energy levels or traps within the bandgap that act as nonradiative recombination centers 7, 8 and reduce carrier lifetimes. E.g., studies have suggested that a Cd vacancy (VCd) in the CdTe lattice could significantly accelerate carrier recombination, leading to a reduction in carrier concentration and a decrease in power conversion efficiency by nearly 5% 9, 10.
CdTe often suffers from low hole density, limiting its efficiency in PV applications 1, 2. To mitigate this, Cu is typically introduced as an acceptor dopant, enhancing hole density 1, 2. Cu doping is typically achieved through high-temperature annealing (673–723 K) in the presence of Cl, followed by a lower-temperature activation step (473–573 K) 11. During CdCl2 treatment, Cl and Cu diffuse into CdTe at concentrations of 1017–1019 cm-3, significantly altering electronic properties 1, 2. While Cui and ClTe act as shallow donors, CuCd behaves as a non-shallow acceptor, even forming defect complexes such as (Cui+CuCd), (Cli+TeCd), and (Cli+CuCd)2+ 12, 13, 14, 15, 16, 3. However, Cu doping typically results in a suboptimal hole density ( 1014 cm-3), far lower than the ideal  1016 cm-3 5.
In contrast, group V doping (mainly As) achieves two to three times higher hole density without sacrificing carrier lifetime or photocurrent 17, 3, 12, 13, 14, 15, 16. Replacing the traditional CdTe absorber layer with Se-alloyed CdSexTe1-x significantly improves CdTe PV efficiency by enhancing long-wavelength absorption, reducing bandgap, increasing short-circuit current density, and improving carrier lifetimes due to better band alignment and reduced interfacial recombination 17, 18, 5. Se passivates grain boundaries and dislocation cores, further extending carrier lifetimes. ZnTe is widely employed as a hole transport layer due to its favorable band alignment with CdTe, facilitating efficient hole collection and enhancing device performance 19. Given these factors, exploring the defect chemical space of Cd/Zn-S/Se/Te bulk and alloyed compositions and interfaces in terms of single native and extrinsic point defects as well as defect complexes, is essential to understanding defect-driven solar cell efficiency and ultimately guiding the development of more efficient and stable CdTe- and CdSeTe-based thin-film PV devices 20.
Defect levels are typically measured experimentally using cathodoluminescence, photoluminescence, optical spectroscopy, or deep-level transient spectroscopy (DLTS) 21. These methods can be challenging due to difficulties in sample preparation and assigning measured levels to specific defects 22. For instance, distinguishing whether an observed peak originates from a particular vacancy, self-interstitial, or an unintended substitutional or interstitial impurity, or a combination of them, is not straightforward 23. To address these challenges, density functional theory (DFT) computations have become a widely used approach for determining the formation energy (Ef) of point defects as a function of Fermi level (EF), the residual charge in the system (q), and chemical potential conditions ($\mu$) 6, 24, 25, 26, 27, 28.
DFT enables the reliable identification of the lowest energy donor- and acceptor-type defects in a solid, the range of possible shallow and deep defect levels, the equilibrium conductivity as determined by the most stable defects (p-type, intrinsic, or n-type), defect concentrations as a function of temperature, electron/hole capture rates, and many other related properties 29, 10, 30, 31, 32, 33, 34, 35, 36. When an appropriate level of theory is applied, DFT-computed charge transition levels often show excellent agreement with experimentally measured levels 29, 23. However, DFT simulations are time-consuming and scale poorly with the number of atoms, causing major issues in using large supercells to represent point defects. It is especially a huge challenge to use DFT for exploring a broad configurational space of potential vacancies, interstitials, antisite defects and defect complexes across dozens of compounds or chemistries of interest 37. Any given defect can adopt numerous configurations and charge states, further adding to the cost of exhaustive exploration 38.
The prediction of defect properties can be greatly accelerated by integrating DFT simulations with state-of-the-art machine learning (ML) approaches such as crystal graph neural networks (GNNs) 39, 40, 41. GNNs have gained immense popularity in recent years and are now widely used to effectively represent and predict energies and properties of molecules, polymers, and solid-state crystalline materials 42, 43. They operate on graph-structured data by transforming crystal structures into crystal graphs where atomic positions are represented as nodes and chemical bonds as edges 44. They are capable of learning intricate internal representations within crystalline environments, which are valuable for predicting various properties of interest such as formation or decomposition energy, electronic bandgap, and defect formation energy, while significantly reducing computational costs.
In past work, we demonstrated the use of GNNs to accelerate the prediction and screening of native defects and functional impurities in the chemical space of group IV, III–V, and II–VI zincblende (ZB) semiconductors 45. This dataset encompasses a wide range of defect types, including vacancies, self-interstitials, anti-site substitutions, and extrinsic interstitial and substitutional defects. While the models were capable of predicting charge-dependent defect formation energies for cation-rich and anion-rich chemical potential conditions directly from an assumed defect structure, they had several limitations: (1) They were trained on a broad chemical space (34 compounds, and practically any element from across the periodic table as a possible defect), which led to large errors on specific compositions and combinations of defects. (2) Models worked well for single native or extrinsic defects in binary compounds, but did not show the same accuracy for ternary or quaternary compositions often encountered in CdTe solar cells (e.g., Se rich CdSexTe1-x local environment and CdxZn1-xTe local environment between CdTe and ZnTe interface), or defect complexes in a variety of compositions 46, 11, 13, 2, 31, 47, 48. (3) They were primarily trained on data from cubic $2\times 2\times 2$ supercells with singular point defects, thus affecting their applicability to larger supercells necessary for defect complexes. (4) They were trained on defect configurations from the cheaper PBE (Perdew–Burke–Ernzerhof) functional within the generalized gradient approximation (GGA), rather than a hybrid functional or other high-fidelity method 49, 50, 51. (5) For obtaining ground state defect configurations, the models must be combined with gradient-free stochastic optimization techniques, preventing them from performing more efficient gradient-based geometry optimization.
Additionally, interfaces—such as those between CdTe (or CdSeTe) and back contact ZnTe—are very important in devices where compositional grading and lattice mismatch can strongly influence dopant incorporation, defect segregation, and carrier recombination 52, 53, 54, 55, 56, 57, 58, 59, 60. Likewise, grain boundaries are inherent to polycrystalline CdTe and often serve as preferential sites for defect clustering, impurity segregation, or charge trapping 61, 62, 63. Understanding point defect energetics in these extended defect structures is therefore critical for improving device stability and performance 64, 61, 65, 62. Our previous defect GNN models did not account for the presence of extended structural features, further underscoring the necessity of extending them to more complex compositions, larger supercells, defect complexes, and higher levels of theoretical accuracy. By incorporating interface and grain boundary configurations into the training data, the model can generalize better to real-world polycrystalline and heterostructured devices, offering predictive insights for defect engineering beyond idealized bulk conditions 66, 67, 68, 69, 70.
The accuracy of ML models is heavily dependent on the quantity and quality of the training data. In the present context, the DFT data for defect structures must encompass all regions of the desired chemical space and all types of defects. By using active learning (AL), new DFT data can be systematically and rationally generated at every step of GNN model retraining by picking new computations that reduce the prediction uncertainty 71, 72, 73, 74, 75, 76, 77. In this work, we used AL to perform new defect calculations to gradually improve the GNN models and ultimately accurately identify low-energy defect configurations in the chemical space of Cd/Zn-Te/Se/S compositions. This specific chemical space is chosen because: (a) Se grading is used in CdTe solar cells, and it is important to predict defect properties as a function of Se concentration, (b) the absorber layer forms interfaces with back contacts such as ZnTe, and Cd-Zn composition grading might thus be important, (c) by tuning the absorber composition itself, defect properties could be manipulated, and (d) examining all low energy defects and dopants across Cd/Zn-Te/Se/S compositions will provide a better library for comparison with experimental measurements, which may arise from different sorts of local coordination environments and chemistries, including defect complexes.
Figure  1: Key challenges limiting CdTe solar-cell performance—point defects, interfaces and grain boundaries, nonradiative carrier recombination, dopant segregation, and the high cost of experiments—and our four-step solution: (i) high-throughput DFT data generation, (ii) machine learning force field (MLFF) training, (iii) rapid large-scale prediction, and (iv) open, public access for defect modeling in CdSexTe1-x solar cells.
While compositions such as ZnS and ZnSe may not directly be of interest to CdTe-based solar cells, they are still chemically informative to the GNN models, and the predicted properties may be useful for understanding local coordination environments that may exist in the solar cell where various components include Cd, Zn, Te, Se, and S. CdS is important because of its use as a buffer layer, though it will likely not be directly relevant as a solar absorber. Interfaces between many of these chalcogenide compounds, which primarily adopt the ZB phase, are certainly relevant to the performance of CdTe and CdSexTe1-x solar cells. Our goal is to develop a model capable of making predictions on suitably large supercells and applicable to defect complexes and grain boundaries in mixed compositions and interfaces. Furthermore, it is vital to extend our predictive capabilities to the non-local hybrid HSE06 (Heyd–Scuseria–Ernzerhof) functional which improves upon the many limitations of semi-local DFT 78.
Our methodology begins with an initial dataset containing bulk and defect configurations in Cd/Zn-Te/Se/S binary and multi-nary compounds, which was used to train early GNN models at GGA-PBE accuracy 79. These models served as the foundation for predicting defect properties across the entire hypothetical defect chemical space that includes thousands of vacancies, interstitials, anti-site defects, substitutional defects, and defect complexes, considering all possible native defects, group V dopants, and unintentional impurities such as Cl and O, which are relevant to CdTe and CdSexTe1-x solar cells 80, 81. To efficiently guide the discovery of new low energy defects, we used Bayesian optimization-based AL, where the next set of DFT simulations was selected based on an acquisition function that aimed to minimize the prediction uncertainty.
Figure 1 captures some of the major challenges in CdTe solar cell performance and our proposed approach for tackling some of them in terms of better defect predictions. Finally, leveraging the augmented and curated HSE06 dataset built upon thousands of bulk and defect configurations optimized with GGA-PBE, we trained a GNN-based machine learning force field (MLFF) model (packaged as “DeFecT-FF") using DFT-derived energies, atomic forces, and stresses. This MLFF is based on the M3GNet framework 44 and once rigorously trained, is capable of optimizing both single defects and defect complexes within the entire Cd/Zn-Te/Se/S chemical space, at both PBE and HSE06 accuracy. The next few sections present details of our methodology and a discussion of the results, ending with some important case studies involving defect formation energy plots from HSE06 with spin-orbit couopling (SOC) that demonstrate the effectiveness and generalizability of the DFT-MLFF approach, and the significant reduction in DFT computational time for defect prediction and screening.
Figure  2: Statistics of the GGA-PBE dataset: (a) Number of bulk configurations corresponding to CdSexTe1-x, CdSxSe1-x, CdxZn1-xS, CdxZn1-xSe, CdxZn1-xTe, Cd0.5Zn0.5SxSe1-x, Cd0.5Zn0.5SexTe1-x, ZnSxSe1-x, and ZnSexTe1-x compositions. (b) Number of neutral defect configurations in CdS, CdSe, CdTe, ZnS, ZnSe, ZnTe, and different CdSexTe1-x compositions, with the inset showing the distribution of vacancy, substitutional, and interstitial defects across the dataset. (c) Bar chart showing the total number of bulk, defect, interface, and grain boundaries (GB) configurations in the dataset. (d) Violin plots showing the distribution of crystal formation energy across all the bulk, defect, interface and grain boundary (GB) structures. Inside each violin, a mini box plot shows the median (central line), quartile, and range (whiskers) (e, f, g) Distribution of defect configurations for different charge states (q = +2, +1, 0, -1, and -2) in bulk defect, interface, and GB configurations, respectively.
acronymacronymacronymacronymacronymacronymacronymacronymacronymacronymacronymacronymacronymacronym Dataset Supercell Size Data Points Bulk dataset from Cd/Zn-Te/Se/S binary, ternary and quaternary alloys $\times2$$\times$2 10080 Bulk dataset from CdSexTe1-x alloys $\times3$$\times$3 26 Defect dataset from 6 Cd/Zn-Te/Se/S binary compounds $\times2$$\times$2 7302 (q=+2), 6201 (q=+1), 8203 (q=0), 6361 (q=-1), 7689 (q=-2) Defect dataset from CdSexTe1-x and CdxZn1-xTe alloys $\times3$$\times$3 10 (q=+2), 12 (q=+1), 26 (q=0), 18 (q=-1), 21 (q=-2) Defect dataset from CdTe/ZnTe interface $\times3$$\times$6 375 (q=+2), 380 (q=+1), 401 (q=0), 381 (q=-1), 330 (q=-2) Defect dataset from CdTe grain boundaries 210 (q=+2), 223 (q=+1), 263 (q=0), 220 (q=-1), 231 (q=-2)
Table 1: Number of data points (or structures) in the GGA-PBE dataset corresponding to different types of bulk or defect configurations, supercell sizes, and charge states.

### v2

@@ Abstract
We developed a framework for predicting the energies and ground state configurations of native point defects, extrinsic dopants and impurities, and defect complexes across zincblende-phase Cd/Zn-Te/Se/S compounds, important for CdTe-based solar cells. This framework, named DeFecT-FF, is powered by high-throughput density functional theory (DFT) computations and crystal graph-based machine learning force field (MLFF) models trained on the DFT data. The Cd/Zn-Te/Se/S chemical space is chosen because alloying at Cd or Te sites is a promising avenue to tailor the electronic and defect properties of the CdTe absorber layer to potentially improve solar cell performance. The sheer number of defect configurations achievable when considering all possible singular defects and their combinations, symmetry-breaking operations, and defect charge states, as well as the expense of running large supercell calculations, makes this an ideal problem for developing accurate and widely-applicable force field models. Here, we introduce our dataset of structures and energies from HSE06 geometry optimization, including bulk and alloyed supercells with and without defects. Data were gradually expanded using active learning and accurate MLFF models were trained to predict energies and atomic forces across different charge states. Via accelerated prediction and screening, we identified many new low energy defect configurations and obtained high-fidelity defect formation energy diagrams using HSE06 calculations with spin-orbit coupling. The DeFecT-FF framework has been released publicly as an online tool on the nanoHUB platform, allowing users to upload any crystallographic information file, generate defects of interest, and compute defect formation energies as a function of Fermi level and chemical potential conditions, thus bypassing expensive DFT calculations.
@@
Md Habibur Rahmana, Maitreyo Biswasa and Arun Mannodi-Kanakkithodia
@@ 1 Introduction
Advancements in solar cell technologies are crucial for meeting global energy demands and supporting the transition to a decarbonized grid Srisuda1, Srisuda2, a2bcx4_abx2, 11133092. Among photovoltaic (PV) technologies, CdTe ranks second after crystalline Si, accounting for about 7% of the global market CdTe_1, Srisuda1, Srisuda2. Its commercial success arises from a direct bandgap of $\sim$1.5 eV, high absorption coefficient ($5\times 10^{5}$ cm-1) CdTe_2, eduardo_men_ndez_proupin_dae1679b, fadhil_k__alfadhili_1e77f8db, o__de_melo_1bf7fe05, kaiying_luo_60e9a185, wei_chao_chen_652bd27c, kai_shen_86d534d1, jinshui_miao_634f66d0, a_gayt_n_garc_a_23ea61bf, xin_zheng_02d025a7, low production cost, and good thin-film conductivity CdTe_3. However, its maximum efficiency of 22.3% remains below the $\sim$30% theoretical limit, mainly due to Shockley-Read-Hall (SRH) recombination associated with grain boundaries, point defects, and dislocations Freysoldt2014-en. Native defects and impurities create trap states within the bandgap that act as nonradiative recombination centers Ganose_Scanlon_2022, Mannodi-Kanakkithodi_2023, such as Cd vacancies (VCd), which accelerate carrier recombination and can reduce power conversion efficiency by nearly 5% sean_1, Turiansky, aneta_wardak_ab3492d5, p__d__hatton_26561f92, michael_a__scarpulla_043ff246.
CdTe often suffers from low hole density, limiting its PV efficiency Srisuda1, Srisuda2. Cu is commonly introduced as an acceptor dopant via high-temperature CdCl2 annealing, where Cu and Cl diffuse at 1017–1019 cm-3, altering electronic properties Yang_Yin, Srisuda1, Srisuda2. While Cui and ClTe act as shallow donors and CuCd as a non-shallow acceptor forming complexes such as (Cui+CuCd) and (Cli+CuCd)2+ Krasikov_1, Krasikov_2, CdTe_1, Cu doping typically yields suboptimal hole density ($\sim$1014 cm-3) compared to the ideal value of $\sim$1016 cm-3 CdTe_3. In contrast, group V dopants such as As achieve higher hole densities without reducing carrier lifetime Ablekim, CdTe_1. Se alloying to create CdSexTe1-x further enhances PV efficiency by improving absorption, band alignment, and carrier lifetimes Ablekim, Fiducia2019-wa. ZnTe, with favorable band alignment, serves as an efficient hole transport layer Gorai_Krasikov. Thus, exploring the defect chemistry of Cd/Zn–S/Se/Te alloys is vital to improving CdTe- and CdSeTe-based thin-film solar cells De_Souza_Harrington_2023.
Defect levels are experimentally measured using cathodoluminescence, photoluminescence, optical spectroscopy, or deep-level transient spectroscopy (DLTS) Lang2003. These methods face significant challenges in sample preparation and assigning measured levels to specific defects Wickramaratne, Kim_Gelczuk. To overcome this, density functional theory (DFT) is widely used to calculate defect formation energy (Ef) as a function of Fermi level (EF), defect charge state (q), and chemical potential ($\mu$) Freysoldt2014-en, Broberg_Bystrom, gw_def, Lee_Din, Grill2004-pk, Buckeridge_2019. DFT enables identification of donor- and acceptor-type defects, shallow or deep defect levels, type of equilibrium conductivity, defect concentrations, and carrier capture rates Mannodi-Kanakkithodi2022-ck, Turiansky, Rahman2023, Defect_Energetics_JPCC, Kim_Park_Hood_Walsh_2019, Lyons2017-sp. When an appropriate level of theory is applied, DFT-computed charge transition levels compare well with experiments Mannodi-Kanakkithodi2022-ck, Kim_Gelczuk, snb_1, snb_2. However, DFT is computationally expensive and scales poorly with system size, making it difficult to explore the vast configurational space of vacancies, interstitials, antisites, and defect complexes across many compounds and charge states mba, Mannodi-HP1.
The prediction of defect properties can be accelerated by integrating DFT simulations with machine learning (ML) approaches such as crystal graph neural networks (GNNs) CGCNN, Choudhary, Chen_Li_Bruna_2017. GNNs effectively represent and predict the energies and properties of molecules, polymers, and crystalline materials Kipf_Welling_2016, Witman_2023 by transforming atomic structures into graphs where atoms are nodes and bonds are edges m3gnet. They learn intricate structural representations to predict properties such as formation or decomposition energy, bandgap, and defect formation energy while reducing computational cost. In prior work, we used GNNs to predict and screen native defects and functional impurities in group IV, III–V, and II–VI zincblende semiconductors Rahman_Gollapalli, covering vacancies, interstitials, anti-site, and extrinsic defects. While the models predicted charge-dependent defect formation energies for different chemical potential conditions, several limitations were observed: (1) training on a broad chemical space (34 compounds) led to large errors for specific compositions; (2) models showed good performance on binaries but lower accuracy for alloy systems such as CdSexTe1-x and CdxZn1-xTe Rahman_Rojsatien, Yang_Yin, Krasikov_2, Srisuda2, Defect_Energetics_JPCC, Rahman_Mannodi-Kanakkithodi_2025b, Lany_soc; (3) reliance on modest 64-atom $2\times 2\times 2$ supercells limited defect complex modeling; (4) use of the semi-local GGA-PBE functional (GGA: Generalized Gradient Approximation, PBE: Perdew–Burke–Ernzerhof) inherently limited prediction fidelity GGA_gaps, HP_hse, Lyons_Van; and (5) dependence on gradient-free optimization with GNN models prevented more efficient gradient-based geometry optimization.
To overcome prior limitations from casting too wide a chemical space, using smaller supercells, and relying on semi-local functionals, we developed a more comprehensive, multi-fidelity methodology. Our approach begins with an initial dataset of bulk and defect configurations spanning Cd/Zn–Te/Se/S binary and multi-nary compounds. These structures were first computed using the PBE functional, providing a baseline set of bulk and defect configurations and their energies, with a substantial portion of the PBE dataset compiled from our previously published works Mannodi-ZB, Mannodi-Kanakkithodi2022-ck. Initial GNN models trained on the PBE dataset  Bapst_Keck served as the foundation for predicting defect properties over a pre-defined defect chemical space containing thousands of vacancies, interstitials, antisites, substitutional defects, and defect complexes. Defect enumeration included all relevant native defects as well as group-V dopants (N, P, As, Sb, Bi) which are promising for achieving p-type conductivity and unintentional impurities such as Cl and O which are known to strongly influence the performance of CdTe and CdSexTe1-x solar cells Li2017-ez, Corsin_2016. The Cd/Zn–Te/Se/S chemical space was chosen due to its relevance to Se grading, Cd–Zn interfaces, and absorber composition tuning in CdTe solar cells, where exploring all low-energy native and extrinsic defects across these compositions provides a comprehensive dataset for experimental comparison. Although ZnS and ZnSe are not primary absorbers, they remain chemically informative, while CdS functions as an important buffer layer CdTe_1, CdTe_2, CdTe_3.
Figure  1: Statistics of the HSE06 dataset: (a) Number of bulk configurations from the CdSexTe1-x, CdSxSe1-x, CdxZn1-xS, CdxZn1-xSe, CdxZn1-xTe, Cd0.5Zn0.5SxSe1-x, Cd0.5Zn0.5SexTe1-x, ZnSxSe1-x, and ZnSexTe1-x compositions. (b) Distribution of defect configurations across the Cd–chalcogen and Zn–chalcogen binaries and ternaries. (c) Violin plots of crystal formation energies (meV/atom) for the entire dataset across five charge states (+2+2 to 2-2). (d,e) Defect formation energy diagrams for CdTe under Cd-rich and Te-rich conditions from HSE06 functional, highlighting the relative stability of key native (VCd, VTe) and extrinsic defects (AsTe, ClTe).
To improve prediction accuracy, active learning was employed to iteratively generate new DFT data and refine the GNN models AL_1, AL_2, AL_3, AL_4, AL_5, Gentle_2010, Farache_Verduzco. Following convergence of the active learning scheme, we performed higher-fidelity HSE06 (Heyd–Scuseria–Ernzerhof)  hse calculations on a curated subset of representative PBE-relaxed structures to obtain more accurate bandgaps, charge transition levels, and defect formation energies. GNN models, specifically using the M3GNet m3gnet architecture for machine learning force fields (MLFFs), were then trained on the HSE06 data and subsequently used for new predictions. Our complete methodology is summarized as: PBE data collection $\rightarrow$ initial PBE GNN model training $\rightarrow$ active-learning–driven expansion of the PBE dataset $\rightarrow$ HSE06 refinement of a subset of the PBE dataset $\rightarrow$ training MLFF models at HSE06 accuracy. The next few sections describe our methodology and results in detail, highlighting the following major contributions of this work:
  • Construction of the largest unified HSE06 defect dataset across Cd/Zn–Te/Se/S compositions, including native and extrinsic defects, and defect complexes, simulated in five charge states.
  • Development of an HSE06 MLFF-based defect optimization workflow that is orders of magnitude faster than full DFT.
  • Release of DeFecT-FF, an online nanoHUB tool defect-ff with the following workflow: input bulk structure + list of defect candidates $\rightarrow$ generation of defect structures with symmetry breaking $\rightarrow$ MLFF optimization across five charge states $\rightarrow$ selection of the lowest-energy configurations $\rightarrow$ final HSE06+SOC (spin–orbit coupling) calculation.